What is the total number of moles present in a 52.0 gram sample of glucose, C6H12O6 (s)? (μc = 12.0 g/mol; μH = 1.0 g/mol; μo = 16.0 g/mol)?


1
Expert's answer
2021-08-19T10:27:03-0400

The molar mass of glucose is:

"M(C_6H_{12}O_6)=6\times{M(C)}+12\times{M(H)}+6\times{M(O)=6\times12.0+12\times1.0+6\times16.0=}180g/mol"


Therefore, the number of moles equals:

"n(C_6H_{12}O_6)=52.0g\times\frac{1mol(C_6H_{12}O_6)}{180g}=0.289mol"


Answer: 0.289 mol